OpenStudy (anonymous):

QRST is a rectangle. If RT= 3X+8 and SQ=X+16, find the value of x

OpenStudy (anonymous):

they both(RT= 3X+8 and SQ=X+16)form the diagonals of a rectangle. and diagonals of a rectangle are equal in length 3x+8=x+16 2x=8 x=4